OpenStudy (anonymous):

The youth group is going on a trip to the state fair. The trip costs $64. Included in that price is $12 for a concert ticket and the cost of 2 passes, one for the rides and one for the game booths. Each of the passes costs the same price. Write an equation representing the cost of the trip, and determine the price of one pass. Solve your equation by showing your work and steps.

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

total price = concert ticket + 2*(cost of pass) find the cost of the pass

OpenStudy (anonymous):

12 + 2 * ( 2 ) ?

Vocaloid (vocaloid):

nope, think carefully total price = 64 concert ticket = 12 total price = concert ticket + 2*(cost of pass) now finish the equation

OpenStudy (anonymous):

64 = 12+2*

OpenStudy (anonymous):

64 = 12 + 2 * ( 2 )

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

also, we can use a variable to represent the cost of the pass, so 64 = 12 + 2x solve for x

OpenStudy (anonymous):

x would equal 26 correct?

OpenStudy (anonymous):

because you have to subtract 12 from 64 which would equals 52 and 26 x 2 equals 52 so x=26

OpenStudy (anonymous):

i got x=26 @Vocaloid

OpenStudy (anonymous):

I think that is correct

Vocaloid (vocaloid):

yes, good

OpenStudy (anonymous):

so complete answer?

OpenStudy (anonymous):

:D

Vocaloid (vocaloid):

geny, read what the question is asking, then read over what we did

OpenStudy (anonymous):

64=12+(2x26) that should be right

OpenStudy (anonymous):

whats the equation? @Vocaloid

OpenStudy (anonymous):

64 = 12 + 2(26)

Vocaloid (vocaloid):

read what I wrote 64 = 12 + 2x...

OpenStudy (anonymous):

the equation is : 64 = 12 + 2x

Vocaloid (vocaloid):

great

OpenStudy (anonymous):

okay thanks :)

OpenStudy (anonymous):

)Fiona has $18 to spend. She spent $4.25, including tax, to buy a notebook. She needs to save $9.75, but she wants to buy a snack. If crackers cost $0.50 per package including tax, what inequality would show the maximum number of packages that Fiona can buy? Solve your inequality by showing your work and steps.

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

0.5x + notebook + amount saved ≤ 18